OpenStudy (anonymous):

Rewrite the formula to find the radius of a sphere. The volume (V) of a sphere is given by the formula V= 4/3 Pi r^3 The sphere’s radius (r) is? a. (3V/4Pi)1/3 B. (4V/3Pi) 1/3 c. (3Pi/4v) 1/3

OpenStudy (haseeb96):

V = 4/3 Pi r^3 3V/4Pi =r^3 r = (3V/4Pi)^1/3

OpenStudy (anonymous):

so its a?

OpenStudy (haseeb96):

yeah